OpenStudy (anonymous):

What is sqrt(38)/2sqrt(6)?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

to rationalize the denominator, you need to multiply top and bottom by sqrt(6)

OpenStudy (anonymous):

???????????????????????????????????????????

jimthompson5910 (jim_thompson5910):

where are you stuck

jimthompson5910 (jim_thompson5910):

no offense, putting up a million question marks doesn't help at all

OpenStudy (anonymous):

srry. I was taught that you don't do anything with the sqrts ?

jimthompson5910 (jim_thompson5910):

the instructions are to rationalize the denominator no?

OpenStudy (anonymous):

it says "What is the exact value of..."

jimthompson5910 (jim_thompson5910):

and the full expression given is \[\large \frac{\sqrt{38}}{2\sqrt{6}}\] right?

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

7.54983443527

jimthompson5910 (jim_thompson5910):

ok you multiply top and bottom by sqrt(6) like this \[\large \frac{\sqrt{38}}{2\sqrt{6}}\] \[\large \frac{\sqrt{38}*\sqrt{6}}{2\sqrt{6}*\sqrt{6}}\] \[\large \frac{\sqrt{38*6}}{2\sqrt{6*6}}\] \[\large \frac{\sqrt{228}}{2\sqrt{36}}\] \[\large \frac{\sqrt{228}}{2*6}\] \[\large \frac{\sqrt{228}}{12}\] \[\large \frac{\sqrt{4*57}}{12}\] \[\large \frac{\sqrt{4}*\sqrt{57}}{12}\] \[\large \frac{2*\sqrt{57}}{12}\] \[\large \frac{\sqrt{57}}{6}\]

OpenStudy (anonymous):

^

OpenStudy (anonymous):

i think i did something wrong no ?

OpenStudy (anonymous):

ok so that is it? Wow...

OpenStudy (anonymous):

THANKYOU!!!!!!!!!!!!!!!

OpenStudy (anonymous):

no problem for nothing lol

jimthompson5910 (jim_thompson5910):

DUDE..IM..A..DUCK you provided the approximate answer they want the exact answer (in terms of radicals)